The lengths of the sides of a triangle are in the extended ratio 4 ​: 9 ​: 10. The perimeter of the triangle is 69 cm. What are the lengths of the​ sides?
Answer by ikleyn(52776)

.

Then the side lengths are 4x, 9x and 10x.


For the perimeter, you have this equation


    4x + 9x + 10x = 69,   or

    23x           = 69,

      x           = 69/23 = 3 cm.


Hence, the sides are 4*3 = 12 cm,  9*3 = 27 cm  and  10*3 = 30 cm.    ANSWER